Staff.

LBTO: Juan Carlos, Brandon

Arcetri:

Summary.

Check Bin1 rotation. yesterday it was not possible. it has been two observing nights after the installation of the clamps.

Run the software check. Brandon to test the pause/resume software.

Close loop with Luci@bin1.

TBD take Luci images and identify some NCPA for different angles.

Bin1 and 4 has new pupil files. The camera lens is inside the range. The close loop for bin1 showed to have some fringing at the edge of the mirror. To be discussed with Enrico. The pause/resume has not been tested yet with the system.

Details.

8:30am telescope released.

Set flat->638 modes.

test preset.

Wed Nov 17 15:46:26.463 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[118] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

Wed Nov 17 15:48:05.791 2021 aos.info - Updated variable AOARB.L.AO_READY: 0

Wed Nov 17 15:48:56.909 2021 aos.PresetAO{FLAO}.complete - PresetAO {FLAO}[118] - dx:nan dy:nan slNull: f1:0.0 f2:1.0 freq:900 mag:0 nBins:1 nModes:501 r0:nan snMode:0 strehl:0 ttMod:0

9:08am

Setting pupil: 20211115_223025

center pupils.

starting point.

The pupils shows a diffx= 1.3 pixels down.

Visually moved down ..

start wobble rotation at this position.

errors goes to 0.63@360

@300,diffx=1.72

@280 diffx=1.7

@260 diffx=1.2

@240 diffx=0.65

@220 diffx=-0.26

Forward.
  • OK
@280, diffx=1.4

@300 diffx= 1.5

@320 diffx=2.0

16:40

loaded pupil #85

x=-1, y=1 from 20211115_223025

@300 diffx=1.0

@280 1.2

forward

@240 diffx=0.2

@280 diffx=0.7

shift x=1, y=2 from 20211115_223025

Pupil:20211117_165356

loaded #86

10:01 AM

10:03 backward. DMs 17:05

@340 diffx=0

forward.

10:11am

good just a little saturaiton in positive X

shift x=1, y=3 from 20211115_223025

pupil:20211117_171232

10:21 am

backward.
  • starting CL 46
  • good
forward

10:33 am

good.

change split from y:4 -> 6,x:4->5

backward.

10:45am

good.

apply shift down in y->5

shift y->5 to get more centered. just to reduce the peak in the positive y

new pupil 20211117_180043

11:08
  • backward
  • CL start y= 22
  • forward
11:17am
  • BACKWARD
  • OK
11:25
  • BACKWARD
  • ok
  • STARTING UP LUCI
  • forward.
  • 11:33am
bin1_20211117.png

bin4

pupil from 20211112_221759

shift x=3, y=0

20211117_194000

split x->2

CL start x=100, y=76

start 12:10PM

backward.

@360 diffx=0 cl->x=100

@220 diffx=0, cly=80

ok for pupil error range.

forward.

12:18PM

@140 cl->x=-34 diffx=0

@260 CL->x=86, diffx=0

@280 CL->x=100, diffx=0.14

@340 CL->x=100,diffx=0.2

shift one pixel in Y down.

x=3, y=1 from 20211112_221759

loaded new.

backward

12:29

@340 CL->x=-4 diffx=0

@260 CL-> x=6, diffx=0

@100 CL->x=-100, dix=-0.2

forward

12:38

@260 CL->x=-18, diffx=0

backward

12:49PM

@100 CL->x=-100 diffx=0

shfit one pixel back in Y.

loaded the pupil 20211117_190400

starting from 0->360

1:02PM

@260 CL->x=97/100 diffx=0.15

@360, CL->x=96,diffx=0.1

bin4-after_clamp-20211117.png

Performance.

2021-11-17 20:24:49.630859| -13006 - Illegal command for state ManualAcquire
2021-11-17 20:25:33.211161| -13006 - Illegal command for state ManualAcquire
  • disturbance1.5
  • bin 1
  • 900hz
  • TN:204847
  • rotation issue?
  • stop preset
  • elab gaves errors that data not available,
Modified AO table to have the pupil for Bin1.

nddir:nan)
Wed Nov 17 20:53:40.751 2021 aos.info - WFS source set to: FLAOWFS
Wed Nov 17 20:53:40.813 2021 aos.info - FLAO: WARNING: added atmospheric disturbance file dist_flao2_KL_v20_atm_s1.5_L040.0_v15.0_ovfreq1000.00_sd3892.fits
Wed Nov 17 20:53:41.688 2021 aos.info - FLAO: WFS: loading setup for LUCIFER
Wed Nov 17 20:53:41.701 2021 aos.info - FLAO: WFS: configuring CCDs

  • disturbance1.5
  • bin 1
  • 900hz
  • TN:20211117_210215
  • rotation issue?
  • TN: _210542
Tracknum 20211117_210542   Closed Loop Yes   Is OK? No - SR may be in error - No compatible luci dark found (i.e. same exposure time or filter or dimensions)   AntiDrift OFF   nphotons/sub/fr -1.000000   Magnitude NaN   # Modes 501   Modal rec Rec_20190930_161000.fits   Binning 1   Frequency [Hz] 900   Modulation 3.000000   FW1 Dichroic 600-1000 nm   FW2 empty   Gain minmax 0.044757 0.845443   Telescope elevation 89.994881   Wind speed 0.000000   number of skipped frames 0   lambda [um] 1.646000   exptime [s] 25.024212   framerate [Hz] 0.039961   no. frames 1   SR SE 0.429970   LUCI dark  
bin1_20211117_s0p8.png

manual gain:

TN:20211117_212229

TN after gopt and optimize

_212643. bad image.

optimize_bin1_20211117.png

connecting with Brandon.

Adsc_stop/start

load the soul latest version...

2:39PM

flat->638 modes

pupil rotator taking

AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Wed Nov 17 22:00:21.915 2021 aos.info - WFS source set to: FLAOWFS
Wed Nov 17 22:01:11.959 2021 aos.PresetAO{FLAO}.failed - PresetAO {FLAO}[127] - [AOException] presetAO: FATAL - exception ::AODefs::AOException{ code = -1 what = [AOException] /home/AOeng/soul/latest/aoroot/calib/wfs/current/ocam2/LUTs/current/pixelGating.txt (code -1010) File open/read/write error File BCUlbt.cpp line 485}

AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Wed Nov 17 22:08:00.650 2021 aos.info - WFS source set to: FLAOWFS

cl,ose loop..

OK

optimize gain:

2021-11-17 22:11:56.475437| -13006 - [AOException] OptimizeGain: FATAL - Cannot detect modal basis loaded on AdSec (-3006) VAR_NOT_FOUND_ERROR

HANDOVER.

3:10PM

-- JuanCarlosGuerra - 17 Nov 2021
Topic revision: r2 - 18 Nov 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback